OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = 10 over quantity x squared minus 7x minus 30 x = 10 and x = 3 x = 10 and x = -3 x = -10 and x = 3 x = -10 and x = -3 Help please im lost

OpenStudy (johnweldon1993):

\[\large \frac{10}{x^2 - 7x - 30}\] Vertical Asymptotes occur when the denominator = 0 So when does \(\large x^2 - 7x - 30\) = 0?

OpenStudy (anonymous):

would it be x=-10 and x=3 ? @johnweldon1993

OpenStudy (johnweldon1993):

Not quite! But close... When you factor \(\large x^2 - 7x - 30\) is is true that you get (x - 10)(x + 3) But when do those = 0? when you plug in x = 10 and x = -3 there...you get 0 So it would be x = 10 and x = -3 Does that make sense?

OpenStudy (anonymous):

yesss makes sense , help me on one more please (: @johnweldon1993

OpenStudy (johnweldon1993):

Sure :)

OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = quantity x minus 4 over quantity x squared plus 13 x plus 36 x = -9 and x = -4 x = -9 and x = 4 x = 9 and x = -4 x = 9 and x = 4 @johnweldon1993

OpenStudy (johnweldon1993):

Alright so same thing here... \[\large \frac{x - 4}{x^2 + 13x + 36}\] So when does that denominator = 0? \[\large x^2 + 13x + 36 = 0\]

OpenStudy (anonymous):

would it be 9 & -4 ? @johnweldon1993

OpenStudy (johnweldon1993):

Hmm, when we factor that...we get what?

OpenStudy (johnweldon1993):

So lets see.... What 2 numbers when multiplied = 36 AND when added = 13? Well 9 times 4 = 36 AND 9 + 4 = 13 so there we go Factored, that would look like (x + 9)(x + 4) We just need to see when that = 0 Well if you have x = -9...you get 0....and when you have x = -4 you also get 0 so x = -9 and x = -4

OpenStudy (anonymous):

thank you so muchhhhhh . will you be on for long its just that im on this chapter that i honestly dont get and its online virtual school , so its not like i have a real teacher & your awesome at explaining this

OpenStudy (johnweldon1993):

Lol sure I'll stay on as long as I can for you :)

OpenStudy (anonymous):

THANKSSSSS (: ! Identify the oblique asymptote of f(x) = quantity x plus 4 over quantity 3 x squared plus 5 x minus 2 y = 3x - 7 y = 3x + 17 y = 0 No oblique asymptote ----------------------------------------- Identify the horizontal asymptote of f(x) = quantity 7 x plus 1 over quantity 2 x minus 9 y = 0 y = negative 1 over 9 y = 7 over 2 No horizontal asymptote -------------------------------------------- Identify the oblique asymptote of f(x) = quantity 2 x squared minus 5 x plus 2 over quantity x minus 3 y = 0 y = 2x + 1 y = 2x - 11 No oblique asymptote --------------------------------------------- Identify the horizontal asymptote of f(x) = 3 over 5 x y = 3 over 5 y = 0 y = 5 over 3 No horizontal asymptote ANY IDEA HOW TO SOLVE THIS :/

OpenStudy (anonymous):

i just tried # 2 would that one be 7 / 2 ? if i did it right......

OpenStudy (johnweldon1993):

Alright Oblique asymptotes are found using long division Vertical asymptotes are found when the denominator = 0 Horizontal asymptotes are found when you see what the function equals when approaching infinity

OpenStudy (johnweldon1993):

So lets do number 1 first :)

OpenStudy (anonymous):

i got No oblique asymptote

OpenStudy (johnweldon1993):

\[\large \frac{x + 4}{3x^2 + 5x - 2}\] And you are correct saying there is NO oblique asymptote :) When there is a higher exponent on the bottom than the top...there will not be an obique asymptote

OpenStudy (johnweldon1993):

Number 2 \[\large \frac{7x + 1}{2x - 9}\] When you set x = infinity...you do INDEED get 7/2 as you showed :) good job!

OpenStudy (johnweldon1993):

What do you get for number 3?

OpenStudy (anonymous):

thanks working #3 now would that one be y=2x+1

OpenStudy (johnweldon1993):

\[\large \frac{2x^2 - 5x + 2}{x - 3} \] So lets see  2x + 1 would be correct! :)

OpenStudy (anonymous):

yayyyyy im getting this & last one i got y=0

OpenStudy (johnweldon1993):

Andddddd that would make you perfect :)

OpenStudy (anonymous):

thanks for the help !

OpenStudy (johnweldon1993):

No problem you did all that work lol :)